OpenStudy (anonymous):

Point (-5,-4) m=undefined into y-y1=m(x-x1) and then into standard form y=mx+b

OpenStudy (azureilai):

(-5,-4)=(x,y) so y-(-5)=m(x-(-4)) y+5=m(x+4)

OpenStudy (anonymous):

and then how would that go into standard form?

OpenStudy (azureilai):

You plug the points you have into the equation -4=m(-5)+b